# DESY-25-144

### v1

@@ Appendix A Maximal transition strength
$\textbf{direct}:$ $\quad\underbrace{\frac{\xi_{w}^{3}}{3}\epsilon}_{\begin{subarray}{c}e_{\mathrm{vac}}/(4\pi)\\ \text{vacuum energy}\phantom{X}\end{subarray}}+\underbrace{\frac{3}{4}\int w_{N}\xi^{2}d\xi}_{\begin{subarray}{c}e_{\mathrm{ini}}/(4\pi)\\ \text{initial thermal energy}\phantom{X}\end{subarray}}=\quad\underbrace{\int\gamma^{2}v^{2}w\xi^{2}d\xi}_{\begin{subarray}{c}e_{K}/(4\pi)\\ \text{fluid motion}\phantom{X}\end{subarray}}\quad+\underbrace{\frac{3}{4}\int w\xi^{2}d\xi}_{\begin{subarray}{c}e_{\mathrm{fin}}/(4\pi)\\ \text{final thermal energy}\phantom{X}\end{subarray}}\ ,$ (76)
$\textbf{inverse}:$ $\quad\underbrace{\frac{3}{4}\int w_{N}\xi^{2}d\xi}_{\begin{subarray}{c}e_{\mathrm{ini}}/(4\pi)\\ \text{initial thermal energy}\phantom{X}\end{subarray}}=\underbrace{\frac{\xi_{w}^{3}}{3}\epsilon}_{\begin{subarray}{c}e_{\mathrm{vac}}/(4\pi)\\ \text{vacuum energy}\phantom{X}\end{subarray}}+\quad\underbrace{\int\gamma^{2}v^{2}w\xi^{2}d\xi}_{\begin{subarray}{c}e_{K}/(4\pi)\\ \text{fluid motion}\phantom{X}\end{subarray}}\quad+\underbrace{\frac{3}{4}\int w\xi^{2}d\xi}_{\begin{subarray}{c}e_{\mathrm{fin}}/(4\pi)\\ \text{final thermal energy}\phantom{X}\end{subarray}}\ .$ (77)
@@
$\textbf{direct}:$ $\quad{\rho_{\rm kin}\over\rho_{\rm tot}}\equiv{\kappa_{\rm direct}\alpha_{N}\over 1+\alpha_{N}}=1-{\int(w/w_{N})\xi^{2}d\xi\over 1+\alpha_{N}}\ \overset{\alpha_{N}\gg 1}{\underset{w/w_{N}\to 0}{\longrightarrow}}1$ (78)
$\textbf{inverse}:$ $\quad{\rho_{\rm kin}\over\rho_{\rm tot}}\equiv{\kappa_{\rm inverse}}=1-|\alpha_{N}|-{\frac{3}{4}\int w\xi^{2}d\xi\over\frac{3}{4}\int w_{N}\xi^{2}d\xi}\penalty 10000\ \penalty 10000\ \overset{\alpha_{N}\to-1/3}{\underset{w/w_{N}\to 0}{\longrightarrow}}\penalty 10000\ \penalty 10000\ 2/3\ ,$ (79)
@@ B.1 Linearised perturbations
Before collisions, each bubble generates a spherically–symmetric self–similar fluid profile of the form
@@
where $\mathbf{r}^{(n)}=\mathbf{x}-\mathbf{x}^{(n)}$ is the radial vector between the $n$-th bubble center and the position $\mathbf{x}$, $T^{(n)}$ is the time since nucleation of the $n$-th bubble, and $v_{\rm ip}(\xi)$ is the invariant velocity profile (the subscript “ip” stands for invariant profile). Because the system is self–similar, the spatial dependence enters only through the ratio $\xi=r^{(n)}/T^{(n)}$.
The Fourier transform of this velocity field is defined as [^8] We are using two different symbols for the velocity: $v$ we refer to the single–bubble velocity, while $u$ is meant for the velocity in the multi–bubble case.
@@
where $v_{L}^{(n)}(q,t)$ is the longitudinal scalar amplitude. To obtain this amplitude explicitly, we can project Eq.˜83 along $\hat{q}_{i}$ and use spherical coordinates with the $\hat{q}$–axis as polar axis. To further make the self–similar scaling explicit, change variables to $\xi=r^{(n)}/T^{(n)}$ and define the dimensionless wave number $z=qT^{(n)}$, therefore
@@ B.2 General definition of the RMS velocity
Relation between $\Omega_{K}$ and the multi–bubble RMS velocity
@@
$\bar{U}_{f,\penalty 10000\ {\rm multi}}^{2}$ $={3\over 4\pi\xi_{w}^{3}}\int_{0}^{\infty}{dz\over(2\pi)^{2}}\,z^{2}\Big[|\phi^{\prime}(z)|^{2}+c_{s}^{2}|\eta(z)|^{2}\Big]$ (104)
$=\bar{U}_{f,\penalty 10000\ {\rm single}}^{2}+{3\over 4\pi\xi_{w}^{3}}\int_{0}^{\infty}{dz\over(2\pi)^{2}}\,z^{2}\,c_{s}^{2}|\eta(z)|^{2}\,.$ (105)
@@
$\kappa_{\rm direct}={e_{K}\over{4\pi\over 3}\xi_{w}^{3}\epsilon},\qquad\kappa_{\rm inverse}={e_{K}\over 4\pi\int_{0}^{\xi_{\rm max}}d\xi\,\xi^{2}\,{3\over 4}\,w_{N}}\,$, (107)
@@
$\Omega_{K}^{\rm direct}$ $=\alpha_{N}\,\kappa_{\rm direct}+{3\Gamma\over 4\pi\xi_{w}^{3}}\int_{0}^{\infty}{dz\over(2\pi)^{2}}\,z^{2}\,c_{s}^{2}|\eta(z)|^{2}\,$, (108)
$\Omega_{K}^{\rm inverse}$ $={3\over 4\pi\xi_{w}^{3}}\,\kappa_{\rm inverse}+{3\Gamma\over 4\pi\xi_{w}^{3}}\int_{0}^{\infty}{dz\over(2\pi)^{2}}\,z^{2}\,c_{s}^{2}|\eta(z)|^{2}\,.$ (109)
@@
Figure 15: Normalized contrast maps of single– vs. multi–bubble efficiencies over $(\xi_{w},\alpha_{N})$, showing $\log\!\big(\Omega_{K}^{\rm single}/\Omega_{K}\big)$. Red regions mark strong discrepancies at low wall velocities, where single–bubble conversion is inefficient while the multi–bubble kinetic fraction remains sizeable due to the accumulation of overlapping acoustic shells.
The relations above make it clear that $\Omega_{K}$ receives two distinct contributions: a first term associated with the single–bubble efficiency (the part directly linked to $\Omega_{K}^{\rm single}$), and a second term emerging from the enthalpy-weighted energy contrast of the overlapping shells. When the wall velocity decreases, the single–bubble contribution vanishes rapidly—this reflects the steep drop of the kinetic energy generated by an isolated expanding wall. In contrast, the enthalpy-weighted term, which originates from the collective overlap of bubbles and from the assumption of linear acoustic evolution after collision, remains finite at low $\xi_{w}$. Intuitively, this can be understood by noting that, for a fixed energy contrast $\alpha_{N}$, the global enthalpy difference between phases persists even if the walls move slowly: the plasma still carries pressure and enthalpy perturbations that continue to drive acoustic motion. Hence, the multi-bubble component encoded in Eq.˜103 dominates at small wall velocities, capturing the sustained coherent motion that the single–bubble efficiency alone cannot describe.

### v2

@@ Appendix A Maximal transition strength
$\textbf{direct}:$ $\quad\underbrace{\frac{\xi_{w}^{3}}{3}\epsilon}_{\begin{subarray}{c}e_{\mathrm{vac}}/(4\pi)\\ \text{vacuum energy}\phantom{X}\end{subarray}}+\underbrace{\frac{3}{4}\int w_{N}\xi^{2}\mathrm{d}\xi}_{\begin{subarray}{c}e_{\mathrm{ini}}/(4\pi)\\ \text{initial thermal energy}\phantom{X}\end{subarray}}=\quad\underbrace{\int\gamma^{2}v^{2}w\xi^{2}\mathrm{d}\xi}_{\begin{subarray}{c}e_{K}/(4\pi)\\ \text{fluid motion}\phantom{X}\end{subarray}}\quad+\underbrace{\frac{3}{4}\int w\xi^{2}\mathrm{d}\xi}_{\begin{subarray}{c}e_{\mathrm{fin}}/(4\pi)\\ \text{final thermal energy}\phantom{X}\end{subarray}}\ ,$ (76)
$\textbf{inverse}:$ $\quad\underbrace{\frac{3}{4}\int w_{N}\xi^{2}\mathrm{d}\xi}_{\begin{subarray}{c}e_{\mathrm{ini}}/(4\pi)\\ \text{initial thermal energy}\phantom{X}\end{subarray}}=\underbrace{\frac{\xi_{w}^{3}}{3}\epsilon}_{\begin{subarray}{c}e_{\mathrm{vac}}/(4\pi)\\ \text{vacuum energy}\phantom{X}\end{subarray}}+\quad\underbrace{\int\gamma^{2}v^{2}w\xi^{2}\mathrm{d}\xi}_{\begin{subarray}{c}e_{K}/(4\pi)\\ \text{fluid motion}\phantom{X}\end{subarray}}\quad+\underbrace{\frac{3}{4}\int w\xi^{2}\mathrm{d}\xi}_{\begin{subarray}{c}e_{\mathrm{fin}}/(4\pi)\\ \text{final thermal energy}\phantom{X}\end{subarray}}\ .$ (77)
@@
$\textbf{direct}:$ $\quad\frac{\rho_{\rm kin}}{\rho_{\rm tot}}\equiv\frac{\kappa_{\rm direct}\alpha_{N}}{1+\alpha_{N}}=1-\frac{\int(w/w_{N})\xi^{2}\mathrm{d}\xi}{1+\alpha_{N}}\ \overset{\alpha_{N}\gg 1}{\underset{w/w_{N}\to 0}{\longrightarrow}}1$ (78)
$\textbf{inverse}:$ $\quad\frac{\rho_{\rm kin}}{\rho_{\rm tot}}\equiv{\kappa_{\rm inverse}}=1-|\alpha_{N}|-\frac{\frac{3}{4}\int w\xi^{2}\mathrm{d}\xi}{\frac{3}{4}\int w_{N}\xi^{2}\mathrm{d}\xi}\penalty 10000\ \penalty 10000\ \overset{\alpha_{N}\to-1/3}{\underset{w/w_{N}\to 0}{\longrightarrow}}\penalty 10000\ \penalty 10000\ 2/3\ ,$ (79)
@@ B.1 Linearised perturbations
Before collisions, each bubble generates a spherically-symmetric self-similar fluid profile of the form
@@
where $\mathbf{r}^{(n)}=\mathbf{x}-\mathbf{x}^{(n)}$ is the radial vector between the $n$-th bubble center and the position $\mathbf{x}$, $T^{(n)}$ is the time since nucleation of the $n$-th bubble, and $v_{\rm ip}(\xi)$ is the invariant velocity profile (the subscript “ip” stands for invariant profile). Because the system is self-similar, the spatial dependence enters only through the ratio $\xi=r^{(n)}/T^{(n)}$.
The Fourier transform of this velocity field is defined as [^7] We are using two different symbols for the velocity: $v$ we refer to the single-bubble velocity, while $u$ is meant for the velocity in the multi-bubble case.
@@
where $v_{L}^{(n)}(q,t)$ is the longitudinal scalar amplitude. To obtain this amplitude explicitly, we can project Eq.˜83 along $\hat{q}_{i}$ and use spherical coordinates with the $\hat{q}-$axis as polar axis. To further make the self-similar scaling explicit, change variables to $\xi=r^{(n)}/T^{(n)}$ and define the dimensionless wave number $z=qT^{(n)}$, therefore
@@ B.2 General definition of the RMS velocity
Relation between $\Omega_{K}$ and the multi-bubble RMS velocity
@@
$\bar{U}_{f,\penalty 10000\ {\rm multi}}^{2}$ $=\frac{3}{4\pi\xi_{w}^{3}}\int_{0}^{\infty}\frac{dz}{(2\pi)^{2}}\,z^{2}\Big[|\phi^{\prime}(z)|^{2}+c_{s}^{2}|\eta(z)|^{2}\Big]$ (104)
$=\bar{U}_{f,\penalty 10000\ {\rm single}}^{2}+\frac{3}{4\pi\xi_{w}^{3}}\int_{0}^{\infty}\frac{dz}{(2\pi)^{2}}\,z^{2}\,c_{s}^{2}|\eta(z)|^{2}\,.$ (105)
@@
$\kappa_{\rm direct}=\frac{e_{K}}{\frac{4\pi}{3}\xi_{w}^{3}\epsilon},\qquad\kappa_{\rm inverse}=\frac{e_{K}}{4\pi\int_{0}^{\xi_{\rm max}}d\xi\,\xi^{2}\,\frac{3}{4}\,w_{N}}\,$, (107)
@@
$\Omega_{K}^{\rm direct}$ $=\alpha_{N}\,\kappa_{\rm direct}+\frac{3\Gamma}{4\pi\xi_{w}^{3}}\int_{0}^{\infty}\frac{dz}{(2\pi)^{2}}\,z^{2}\,c_{s}^{2}|\eta(z)|^{2}\,$, (108)
$\Omega_{K}^{\rm inverse}$ $=\frac{3}{4\pi\xi_{w}^{3}}\,\kappa_{\rm inverse}+\frac{3\Gamma}{4\pi\xi_{w}^{3}}\int_{0}^{\infty}\frac{dz}{(2\pi)^{2}}\,z^{2}\,c_{s}^{2}|\eta(z)|^{2}\,.$ (109)
@@
Figure 15: Normalized contrast maps of single- vs. multi-bubble efficiencies over $(\xi_{w},\alpha_{N})$, showing $\log\!\big(\Omega_{K}^{\rm single}/\Omega_{K}\big)$. Red regions mark strong discrepancies at low wall velocities, where single-bubble conversion is inefficient while the multi-bubble kinetic fraction remains sizeable due to the accumulation of overlapping acoustic shells.
The relations above make it clear that $\Omega_{K}$ receives two distinct contributions: a first term associated with the single-bubble efficiency (the part directly linked to $\Omega_{K}^{\rm single}$), and a second term emerging from the enthalpy-weighted energy contrast of the overlapping shells. When the wall velocity decreases, the single-bubble contribution vanishes rapidly—this reflects the steep drop of the kinetic energy generated by an isolated expanding wall. In contrast, the enthalpy-weighted term, which originates from the collective overlap of bubbles and from the assumption of linear acoustic evolution after collision, remains finite at low $\xi_{w}$. Intuitively, this can be understood by noting that, for a fixed energy contrast $\alpha_{N}$, the global enthalpy difference between phases persists even if the walls move slowly: the plasma still carries pressure and enthalpy perturbations that continue to drive acoustic motion. Hence, the multi-bubble component encoded in Eq.˜103 dominates at small wall velocities, capturing the sustained coherent motion that the single-bubble efficiency alone cannot describe.